40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 2 21:16:32 2009, Yoichi, Update, Locking, Low-gain High-gain PD switching may not be working well 
    Reply  Tue Mar 3 19:26:44 2009, Yoichi, Update, Locking, Low-gain High-gain PD switching may not be working well 
       Reply  Tue Mar 3 23:59:26 2009, Yoichi, Update, Locking, Low-gain High-gain PD switching may not be working well 
          Reply  Wed Mar 4 03:37:51 2009, Yoichi, Update, Locking, Low-gain High-gain PD switching may not be working well 
Message ID: 1351     Entry time: Tue Mar 3 23:59:26 2009     In reply to: 1350     Reply to this: 1352
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Low-gain High-gain PD switching may not be working well 

Quote:
I checked the switching of the QPDX from high gain to low gain.
Switching happens as expected, but the low gain QPDX output was very low compared to QPDY.
Also the digital gain for the high gain TRX was not matched with the low gain one. So when the switching happens, there is a large jump in the TRX.

I also found that the offset values for the low gain QPDX were totally wrong. I adjusted it.
Then I removed a beam splitter in front of the QPDX to increase the power falling on it.
But still the low gain QPDX output is four times lower than the low gain QPDY.

I'm still working on it. So don't expect the switching to work correctly at this moment.
I'm planning to be back after the dinner.



This sounds like the QPD whitening gain sliders may be stuck. The slider twiddling script should be run, or the sliders should be twiddled by hand.
ELOG V3.1.3-